OpenStudy (anonymous):

i cant seem to find the derivitave of the square root(3x+1). Any idea how to help?

OpenStudy (anonymous):

d/dx (3x+1)^(1/2) = (1/2)(3x+1)^(-1/2) (3) = (3/2)(3x+1)^(-1/2)

OpenStudy (anonymous):

\[\frac{3}{2 \sqrt{3 x+1}}\] Just use the chain rule. The inside or u would be 3x+1

OpenStudy (anonymous):

SUPER THANKS!! We havent been taught this chain rule tho i have seen it talked about on google. I came up with 3/3x+1

OpenStudy (anonymous):

um... d/dx(3x+1)^(1/2) to change the square root into a power. for the power rule...

OpenStudy (anonymous):

the answer is (1/2)*(3x+1)^(1/2)*(3), you can learn more about the power rule here: http://en.wikipedia.org/wiki/Calculus_with_polynomials#Power_rule